OpenStudy (anonymous):

MEDAL GIVEN! Rewrite each expression in the form ax^n. Remember:(ax^m)^n=a^n x^mn a. (16x^28)^-5/4 b. (-32x^-30)^-6/5

OpenStudy (anonymous):

a is the coefficient, x is a variable and we know that n is the power. Looking at the equation a = 16 n = -5/4 put it in the form ax^n a. ax^n = 16x^-5/4 can you solve for b.?

OpenStudy (anonymous):

so i did it this method.. (16)^-5/4 and x^28^-5/4 since you have to multiply exponents 28 and -5/4 for the x the product would be x^-35 as for the 4√16^-5 i'm not sure how you would write that

OpenStudy (anonymous):

would you do 2x2x2x2 which is 16 2^4 but 2^-5 isn't a clean value its a decimal

OpenStudy (anonymous):

i get .03125 times x^-35 as my final answer but i'm not really positive its correct

OpenStudy (anonymous):

yea 2^4 = 16. My question is once you get it in the form do you have to solve for the equation (ax^m)^n=a^n x^mn? Since the only thing i see it asks is rewrite the expression in the form ax^n. That technically doesn't mean that you have to solve for it Yea but you are doing it right

OpenStudy (anonymous):

they just want you to keep it in that form and solve the value in the square root.. the reason why I wasn't sure if I did the math correct was because the 4√16^-5 was raised to a negative exponent which threw me off a little

OpenStudy (anonymous):

no its correct as long as the coefficient isn't a negative number, which if it was would leave it as an imaginary number

OpenStudy (anonymous):

oh ok.. because for some reason i thought maybe it would be 1/16^-5? lol and yea that makes sense because it's asking for it to be in the for ax^n

OpenStudy (anonymous):

hmm i do this both ways let me see what i get but the result will be postive with a negative power

OpenStudy (anonymous):

the 2 raised to the -5

OpenStudy (dan815):